# Towards Principled Multi-Agent Task Agnostic Exploration

### v1

@@ 4 A Formal Characterization of Multi-Agent Task-Agnostic Exploration
In the previous section, we described how different objectives enforce different behaviors for task-agnostic explorative policies. In this section, we address the second research question:
(ii) Are different formulations related in some way and when crucial differences emerge?
First of all, we show that if we look at task-agnostic exploration tasks, i.e. the ones defined by setting the functional $\mathcal{F}(\cdot):=H(\cdot)$, all the objectives in infinite-trials formulation can be elegantly linked one to the other though the following result:
@@
For every Convex Markov Game $\mathcal{M}_{H}$ equipped with an entropy functional, for a fixed (joint) policy $\pi=(\pi^{i})_{i\in\mathcal{N}}$ the infinite-trials objectives are ordered according to:
$\frac{H(d^{\pi})}{|\mathcal{N}|}$ $\leq\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}H(d_{i}^{\pi})\leq H(\tilde{d}^{\pi})$
$H(\tilde{d}^{\pi})\leq\sup_{i\in[\mathcal{N}]}$ $H(d_{i}^{\pi})+\log(|\mathcal{N}|)\leq H(d^{\pi})+\log(|\mathcal{N}|)$
The full derivation of these bounds is reported in Appendix A. This set of bounds prescribe that the difference in performances over infinite-trials objective for the same policy can be generally bounded as a function of the number of agents. In particular, disjoint objectives generally provides poor approximations of the joint objective from the point of view of the single-agent, while the mixture objective is guaranteed to be a rather good lower bound to the joint entropy as well, since its over-estimation scales logarithmically with the number of agents.
It is still an open question how hard it is to actually optimize for these objectives. Now, while CMGs are a novel interaction framework, whose general properties are far from being well-understood, they surely enjoy some nice properties. In particular, as commonly done in Potential Markov Games (Leonardos et al., 2021), it is possible to exploit the fact that performing Policy Gradient (PG, Sutton et al., 1999; Peters & Schaal, 2008) independently among the agents is equivalent to running PG jointly, when this is done over the same common objective (Appendix A.1, Lemma A.5). This allows us to provide a rather positive answer, here stated informally and extensively discussed in Appendix A.1 :
@@
Under proper assumptions, for every CMG $\mathcal{M}_{\mathcal{F}}$, independent Policy Gradient over infinite trials non-disjoint objectives via centralized policies of the form $\pi=(\pi^{i}\in\Delta_{\mathcal{S}}^{\mathcal{A}^{i}})_{i\in[\mathcal{N}]}$ converges fast.
@@
This result suggests that PG should be generally enough for the infinite-trials optimization, and thus, from a certain point of view, these problems might not be of so much interest. However, convex MDP theory has outlined that optimizing for infinite-trials objectives might actually lead to extremely poor performances as soon as the policies are deployed over just a handful of trials, i.e. in almost any practical scenario (Mutti et al., 2023a). We show that this property transfers almost seamlessly to CMGs as well, with interesting additional take-outs:
Theorem 4.2 (Objectives Mismatch in CMGs).
@@
For every CMG $\mathcal{M}_{\mathcal{F}}$ equipped with a $L$-Lipschitz function $\mathcal{F}$, let $K\in\mathbb{N}^{+}$ be a number of evaluation episodes/trials, and let $\delta\in(0,1]$ be a confidence level, then for any (joint) policy $\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}$, it holds that
$|\zeta_{K}(\pi)-\zeta_{\infty}(\pi)|\leq LT\sqrt{\frac{2|\mathcal{S}|\log(2T/\delta)}{K}},$
$|\zeta^{i}_{K}(\pi)-\zeta^{i}_{\infty}(\pi)|\leq LT\sqrt{\frac{2|\tilde{\mathcal{S}}|\log(2T/\delta)}{K}},$
$|\tilde{\zeta}_{K}(\pi)-\tilde{\zeta}_{\infty}(\pi)|\leq LT\sqrt{\frac{2|\tilde{\mathcal{S}}|\log(2T/\delta)}{|\mathcal{N}|K}}.$
In general, this set of bounds confirms that infinite and finite trials objectives might be extremely different, and thus optimizing the infinite-trials objective might lead to unpredictable performance at deployment, whenever this is done over a handful of trials. This property is inherently linked to the convex nature of convex MDPs, and Mutti et al. (2023a) introduces it to highlight that the concentration properties of empirical state-distributions (Weissman et al., 2003) allow for a nice dependency on the number of trials in controlling the mismatch. In multi-agent settings, the result portraits a more nuanced scene:
@@
(i) The mismatch still scales with the cardinality of the support of the state distribution, yet, for joint objectives, this quantity scales very poorly in the number of agents. [^4] Indeed, in the case of product state-spaces $\mathcal{S}=\times_{i\in[\mathcal{N}]}\mathcal{S}_{i}$ the cardinality scales exponentially with the number of agents $|\mathcal{N}|$ Thus, even though optimizing infinite-trials joint objectives might be rather easy in theory as Fact 4.1 suggests, it might result in poor performances in practice. On the other hand, the quantity is independent of the number of agents for disjoint and mixture objectives.
(ii) Looking at mixture objectives, the mismatch scales sub-linearly with the number of agents $\mathcal{N}$. Thus, in some sense, the number of agents has the same role as the number of trials: the more the agents the less the deployment mismatch, and at the limit, with $\mathcal{N}\rightarrow\infty$, the mismatch vanishes completely. [^5] One should note that in this scenario, though, all the bounds of Lemma 4.1 linking different objectives become vacuous. In other words, this result portraits a striking difference with respect to joint objectives: when facing task-agnostic exploration over mixtures, a reasonably high number of agents compared to the size of the state-space actually helps, and simple policy gradient over mixture objectives might be enough.
Remarks.
One should notice that the results of Fact 4.1 are valid only for specific classes of policies, namely centralized policies of the form $\pi=(\pi^{i}\in\Delta_{\mathcal{S}}^{\mathcal{A}^{i}})_{i\in[\mathcal{N}]}$. Up to our knowledge, no guarantees are known for decentralized policies even in linear MGs. Interestingly though, the finite-trials formulation do offer additional insights on the behavior of optimal decentralized policies, a striking difference with respect to both the infinite-trial objectives and the linear MG interaction model in general. The interested reader can learn more about this in Appendix A.2.
5 Trust Region for Exploration in Practice
@@
As stated before, a core drive of this work is addressing multi-agent task-agnostic exploration in practical scenarios. Yet, these cases are also the ones in which performing PG of infinite-trials objectives provide poor performance guarantees at deployment. In other words, here we address the third research question, that is:
(iii) How can we explicitly address multi-agent task-agnostic exploration in practical scenarios?
@@
To do so, our attention will focus on the finite trials objectives explicitly, more specifically on the single-trial case with $K=1$. Remarkably, it is possible to directly optimize the single-trial objective in multi-agent cases with decentralized algorithms: we introduce Trust Region Pure Exploration (TRPE), the first decentralized algorithm that explicitly addresses single-trial objectives in CMGs, with task-agnostic exploration as a special case. TRPE takes inspiration from trust-region based methods as TRPO (Schulman et al., 2017), as they recently enjoyed an ubiquitous success and interest for their surprising effectiveness in multi-agent problems (Yu et al., 2022).
In fact, trust-region analysis nicely align with the properties of finite-trials formulations and allow for an elegant extension to CMGs through the following.
@@
For every CMG $\mathcal{M}_{\mathcal{F}}$ equipped with a $L$-Lipschitz function $\mathcal{F}$, let $d_{1}$ be a general single-trial distribution $d_{1}=\{d_{1},d_{1,i},\tilde{d}_{1}\}$, then for any per-agent deviation over policies $\pi=(\pi^{i},\pi^{-i})$, $\tilde{\pi}=(\tilde{\pi}^{i},\pi^{-i})$, it is possible to define a per-agent Surrogate Function $\mathcal{L}^{i}(\tilde{\pi}/\pi)$ of the form
$\mathcal{L}^{i}(\tilde{\pi}/\pi)=\operatorname*{\mathbb{E}}_{d_{1}\sim p^{\pi}_{1}}\rho^{i}_{\tilde{\pi}/\pi}\mathcal{F}(d_{1}),\vspace{-0.2cm}$
where $\rho^{i}$ is the per-agent importance-weight coefficient $\rho^{i}_{\tilde{\pi}/\pi}=p^{\tilde{\pi}}_{1}/p^{\pi}_{1}=\prod_{t\in[T]}\frac{\tilde{\pi}^{i}(\mathbf{a}^{i}[t]|\mathbf{s}^{i}[t])}{\pi^{i}(\mathbf{a}^{i}[t]|\mathbf{s}^{i}[t])}$, such that for $\zeta_{1}\in\{\zeta_{1}^{\infty},\zeta_{1}^{i},\tilde{\zeta}_{1}\}$.
Algorithm 1 Trust Region Pure Exploration (TRPE)
  Input: exploration horizon $T$, number of trajectories $N$, trust-region threshold $\delta$, learning rate $\eta$.
  initialize $\bm{\theta}=(\theta^{i})_{i\in[\mathcal{N}]}$
  for epoch = $1,2,\ldots,$ until convergence do
     Collect $N$ trajectories with $\pi_{\bm{\theta}}=(\pi^{i}_{\theta^{i}})_{i\in[\mathcal{N}]}$.
     for agent $i=1,2,\ldots,$ concurrently do
        Construct datasets $\mathcal{D}^{i}=\{(\mathbf{s}^{i}_{n},\mathbf{a}^{i}_{n}),\zeta_{1}^{n}\}_{n\in[N]}$
        $\theta^{i}\leftarrow\emph{IS-Optimizer}(\mathcal{D}^{i},\theta^{i})$
     end for
  end for
  Output: task-agnostic exploration (joint) policy $\pi_{\bm{\theta}}=(\pi^{i}_{\theta^{i}})_{i\in[\mathcal{N}]}$
IS-Optimizer
  Input: Dataset $\mathcal{D}^{i}$, sampling parameter $\theta^{i}$.
  Initialize $h=0$ and $\theta^{i}_{h}=\theta^{i}$
  while  $D_{\mathrm{KL}}(\pi^{i}_{\theta^{i}_{h}}\|\pi^{i}_{\theta^{i}_{0}})\leq\delta$ do
     Compute ^$\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$ via IS
     Perform Gradient step $\theta^{i}_{h+1}=\theta^{i}_{h}+\eta\nabla_{\theta^{i}_{h}}\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$
     $h\leftarrow h+1$
  end while
  Output: parameters $\bm{\theta}_{h}$
@@
From this definition, it follows that the trust-region algorithmic blueprint of Schulman et al. (2017) can be directly applied to single-trial formulations, with per-agent policies within a parametric space of stochastic differentiable policies $\Theta=\{\pi^{i}_{\theta^{i}}:\theta^{i}\in\Theta^{i}\subseteq\mathbb{R}^{q}\}$. In practice, kl-divergence is employed for greater scalability provided a trust-region threshold $\delta$, we address the following optimization problem for each agent:
$\max_{\tilde{\theta}^{i}\in\Theta^{i}}\mathcal{L}^{i}(\tilde{\theta}^{i}/\theta^{i}),$

### v3

@@ 4 A Formal Characterization of Multi-Agent task-agnostic exploration
In the previous section, we provided a principled problem formulation of multi-agent task-agnostic exploration through an array of different objectives. Here, we address the second research question:
(Q2) How are different formulations related? Do crucial theoretical differences emerge?
First of all, we show that all the entropy infinite-trials objectives can be linked through the following:
@@
For every cMG $\mathcal{M}_{H}$, for a fixed (joint) policy $\pi=(\pi^{i})_{i\in\mathcal{N}}$ the infinite-trials objectives are ordered according to:
$\frac{H(d^{\pi})}{|\mathcal{N}|}\leq\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}H(d_{i}^{\pi})\leq H(\tilde{d}^{\pi})\leq\sup_{i\in[\mathcal{N}]}H(d_{i}^{\pi})+\log(|\mathcal{N}|)\leq H(d^{\pi})+\log(|\mathcal{N}|)$
The full derivation of these bounds is reported in Appendix B. This set of bounds characterizes the performances over infinite-trials objective for the same policy as a function of the number of agents. In particular, disjoint objectives generally provide poor approximations of the joint objective, while the mixture objective is guaranteed to be a rather good lower bound to the joint entropy as well, since its over-estimation scales logarithmically with the number of agents.
It is still an open question how hard it is to actually optimize for these objectives. Here, we provide a rather positive answer, here stated informally and extensively discussed in Appendix B.1:
@@
Under proper assumptions, for every cMG $\mathcal{M}_{\mathcal{F}}$, independent Policy Gradient over infinite trials non-disjoint objectives (see Eqs. (1),(3)) via centralized-information policies of the form $\pi=(\pi^{i}\in\Delta_{\mathcal{S}}^{\mathcal{A}^{i}})_{i\in[\mathcal{N}]}$ converges fast.
@@
This result suggests that PG should be generally enough for the infinite-trials optimization. However, cMDP theory has outlined that optimizing for infinite-trials objectives might actually lead to extremely poor performance as soon as the policies are deployed over just a handful of trials, i.e. in almost any practical scenario [29]. We show that this property transfers to cMGs, with interesting additional takeouts:
Theorem 4.2 (Finite-Trials Mismatch in cMGs).
@@
For every cMG $\mathcal{M}_{\mathcal{F}}$ equipped with a $L$-Lipschitz function $\mathcal{F}$, let $K\in\mathbb{N}^{+}$ be a number of evaluation episodes/trials, and let $\delta\in(0,1]$ be a confidence level, then for any (joint) policy $\pi=(\pi^{i}\in\Pi^{i})_{i\in[\mathcal{N}]}$, it holds that the mismatches $err(\pi):=|\zeta_{K}(\pi)-\zeta_{\infty}(\pi)|,err^{i}(\pi):=|\zeta^{i}_{K}(\pi)-\zeta^{i}_{\infty}(\pi)|,\tilde{err}(\pi):=|\tilde{\zeta}_{K}(\pi)-\tilde{\zeta}_{\infty}(\pi)|$ are bounded as:
$err(\pi)\leq LT\sqrt{\frac{2|\mathcal{S}|\log(2T/\delta)}{K}},\quad err^{i}(\pi)\leq LT\sqrt{\frac{2|\tilde{\mathcal{S}}|\log(2T/\delta)}{K}},\quad\tilde{err}(\pi)\leq LT\sqrt{\frac{2|\tilde{\mathcal{S}}|\log(2T/\delta)}{|\mathcal{N}|K}}.$
@@
In general, this set of bounds confirms that for any given policy, infinite and finite trials performances might be extremely different, and thus optimizing the infinite-trials objective might lead to unpredictable performance at deployment, whenever this is done over a handful of trials. This property is inherently linked to the convex nature of cMGs, and compared to cMDPs [29], in multi-agent settings the result portraits a more nuanced scene: (i) The mismatch still scales with the cardinality of the support of the state distribution, yet, for joint objectives, this quantity scales very poorly in the number of agents. [^4] Indeed, in the case of product state-spaces $\mathcal{S}=\times_{i\in[\mathcal{N}]}\mathcal{S}_{i}$ the cardinality scales exponentially with the number of agents $|\mathcal{N}|$. Thus, even though optimizing infinite-trials joint objectives might be rather easy in theory as Fact 4.1 suggests, it might result in poor performances in practice. On the other hand, the quantity is independent of the number of agents for disjoint and mixture objectives. (ii) Looking at mixture objectives, the mismatch scales sub-linearly with the number of agents $\mathcal{N}$. In some sense, the number of agents has the same role as the number of trials: The more the agents the less the deployment mismatch, and at the limit, with $\mathcal{N}\rightarrow\infty$, the mismatch vanishes completely. [^5] In this scenario, all the bounds of Lemma 4.1 linking different objectives become vacuous.
5 An Algorithm for Multi-Agent task-agnostic exploration in Practice
@@
As stated before, a core drive of this work is addressing practical scenarios, where only a handful of trials can be drawn while interacting with the environment. Yet, Th. 4.2 implies that optimizing for infinite-trials objectives, as with PG updates in Fact 4.1, might result in poor performance at deployment. As a result, here we address the third research question, that is:
(Q3) Can we explicitly pre-train a policy for task-agnostic exploration
in practical multi-agent scenarios?
@@
Remarkably, it is possible to directly optimize the single-trial objective in multi-agent cases with decentralized algorithms: We introduce Trust Region Pure Exploration (TRPE), the first decentralized algorithm that explicitly addresses single-trial objectives in cMGs, with task-agnostic exploration as a special case. TRPE takes inspiration from trust-region based methods as TRPO [36].
@@
For every cMG $\mathcal{M}_{\mathcal{F}}$ equipped with a $L$-Lipschitz function $\mathcal{F}$, let $d_{1}$ be a general single-trial distribution $d_{1}=\{d_{1},d_{1,i},\tilde{d}_{1}\}$, then for any per-agent deviation over policies $\pi=(\pi^{i},\pi^{-i})$, $\tilde{\pi}=(\tilde{\pi}^{i},\pi^{-i})$, it is possible to define a per-agent Surrogate Function $\mathcal{L}^{i}(\tilde{\pi}/\pi)$ of the form $\mathcal{L}^{i}(\tilde{\pi}/\pi)=\operatorname*{\mathbb{E}}_{d_{1}\sim p^{\pi}_{1}}\rho^{i}_{\tilde{\pi}/\pi}\mathcal{F}(d_{1}),$ where $\rho^{i}$ is the per-agent importance-weight coefficient $\rho^{i}_{\tilde{\pi}/\pi}=p^{\tilde{\pi}}_{1}/p^{\pi}_{1}=\prod_{t\in[T]}\frac{\tilde{\pi}^{i}(\mathbf{a}^{i}[t]|\mathbf{s}^{i}[t])}{\pi^{i}(\mathbf{a}^{i}[t]|\mathbf{s}^{i}[t])}$.
Algorithm: Trust Region Pure Exploration (TRPE) 1:  Input: exploration horizon $T$, trajectories $N$, trust-region threshold $\delta$, learning rate $\eta$ 2:  Initialize $\bm{\theta}=(\theta^{i})_{i\in[\mathcal{N}]}$ 3:  for epoch = $1,2,\ldots$ until convergence do 4:     Collect $N$ trajectories with $\pi_{\bm{\theta}}=(\pi^{i}_{\theta^{i}})_{i\in[\mathcal{N}]}$ 5:     for agent $i=1,2,\ldots$ concurrently do 6:        Set datasets $\mathcal{D}^{i}=\{(\mathbf{s}^{i}_{n},\mathbf{a}^{i}_{n}),\zeta_{1}^{n}\}_{n\in[N]}$ 7:        $h=0$, $\theta^{i}_{h}=\theta^{i}$ 8:        while $D_{\mathrm{KL}}(\pi^{i}_{\theta^{i}_{h}}\|\pi^{i}_{\theta^{i}_{0}})\leq\delta$ do 9:           Compute ^$\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$ via IS. 10:           $\theta^{i}_{h+1}=\theta^{i}_{h}+\eta\nabla_{\theta^{i}_{h}}\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$ 11:           $h\leftarrow h+1$ 12:        end while 13:        $\theta^{i}\leftarrow\theta^{i}_{h}$ 14:     end for 15:  end for 16:  Output: joint policy $\pi_{\bm{\theta}}=(\pi^{i}_{\theta^{i}})_{i\in[\mathcal{N}]}$
@@
From this definition, it follows that the trust-region algorithmic blueprint of [36] can be directly applied to single-trial formulations, with a parametric space of stochastic differentiable policies for each agent $\Theta=\{\pi^{i}_{\theta^{i}}:\theta^{i}\in\Theta^{i}\subseteq\mathbb{R}^{q}\}$.
In practice, KL-divergence is employed for greater scalability provided a trust-region threshold $\delta$, we address the following optimization problem for each agent:
$\max_{\tilde{\theta}^{i}\in\Theta^{i}}\mathcal{L}^{i}(\tilde{\theta}^{i}/\theta^{i})\;\text{s.t. }D_{\mathrm{KL}}(\pi^{i}_{\tilde{\theta}^{i}}\|\pi^{i}_{\theta^{i}})\leq\delta$
The main idea then follows from noticing that the surrogate function in Def. 5.1 consists of an Importance Sampling (IS) estimator [30], and it is then possible to optimize it in a fully decentralized and off-policy manner [22, 24]. More specifically, given a pre-specified objective of interest $\zeta_{1}\in\{\zeta_{1}^{\infty},\zeta_{1}^{i},\tilde{\zeta}_{1}\}$, agents sample $N$ trajectories $\{(\mathbf{s}_{n},\mathbf{a}_{n})\}_{n\in[N]}$ following a (joint) policy with parameters $\bm{\theta}_{0}=(\theta^{i}_{0},\theta^{-i}_{0})$. They then compute the values of the objective for each trajectory, building separate datasets $\mathcal{D}^{i}=\{(\mathbf{s}^{i}_{n},\mathbf{a}^{i}_{n}),\zeta_{1}^{n}\}_{n\in[N]}$ and using it to compute the Monte-Carlo approximation of the Surrogate Function, namely ^$\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})=\frac{1}{N}\sum_{n\in[N]}\rho^{i,n}_{\theta^{i}_{h}/\theta^{i}_{0}}\zeta^{n}_{1},\quad\rho^{i,n}_{\theta^{i}_{h}/\theta^{i}_{0}}=\prod_{t\in[T]}\pi^{i}_{\theta^{i}_{h}}(\mathbf{a}^{i}_{n}[t]|\mathbf{s}^{i}_{n}[t])/\pi^{i}_{\theta^{i}_{0}}(\mathbf{a}^{i}_{n}[t]|\mathbf{s}^{i}_{n}[t]),$ and $\zeta^{n}_{1}$ is the plug-in estimator of the entropy based on the empirical measure $d_{1}$ [31]. Finally, at each off-policy iteration $h$, each agent updates its parameter via gradient ascent $\theta^{i}_{h+1}\leftarrow\theta^{i}_{h}+\eta\nabla_{\theta^{i}_{h}}\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$ until the trust-region boundary is reached, i.e., when it holds $D_{\mathrm{KL}}(\pi^{i}_{\tilde{\theta}^{i}}\|\pi^{i}_{\theta^{i}})>\delta.$ The pseudo-code of TRPE is reported in Algorithm 5. We remark that even though TRPE is applied to task-agnostic exploration, the algorithmic blueprint does not explicitly require the function $\mathcal{F}$ to be the entropy function and thus it is of independent interest. The interested reader can found some comments of its potential limitations in Appendix C.
6 Empirical Corroboration
In this section, we address the last research question, that is:
(Q4) Do crucial differences emerge in practice? Does this have
an impact on downstream tasks learning?
by providing empirical corroboration of the findings discussed so far. Especially, we aim to answer the following questions: (Q4.1) Is Algorithm 5 actually capable of optimizing finite-trials objectives? (Q4.2) Do different objectives enforce different behaviors, as expected from Section 3? (Q4.3) Does the clustering behavior of mixture objectives play a crucial role? If yes, when and why?
Experimental Domains.  The experiments were performed with the aim to illustrate essential features of task-agnostic exploration suggested by the theoretical analysis, and for this reason the domains were selected for being challenging while keeping high interpretability. The first is a notoriously difficult multi-agent exploration task called secret room [MPE, 20], [^6] We highlight that all previous efforts in this task employed centralized-information policies. On the other hand, we are interested on the role of the entropic feedback in fostering coordination rather than full-state conditioning, thus we employed decentralized-information policies. referred to as Env. (i). In such task, two agents are required to reach a target while navigating over two rooms divided by a door. In order to keep the door open, at least one agent have to remain on a switch. Two switches are located at the corners of the two rooms. The hardness of the task then comes from the need of coordinated exploration, where one agent allows for the exploration of the other. The second is a simpler exploration task yet over a high dimensional state-space, namely a 2-agent instantiation of Reacher [MaMuJoCo, 32], referred to as Env. (ii). Each agent corresponds to one joint and equipped with decentralized-information policies. In order to allow for the use of plug-in estimator of the entropy [31], each state dimension was discretized over 10 bins.
MixtureJointDisjointUniform
(a) Joint Entropy
(b) Mixture Entropy
Figure 1: Single-trial Joint and Mixture Entropy induced by different objective optimization along a $T=50$ horizon. (Right) State Distributions of two agents induced by different learned policies. We report the average and 95% confidence interval over 4 runs.
Refer to caption (c) Mixture Refer to caption (d) Joint Refer to caption (e) Disjoint Refer to caption (f) Uniform
task-agnostic exploration.  As common for the unsupervised RL framework [8, 13, 18, 25], Algorithm 5 was first tested in her ability to optimize for task-agnostic exploration objectives, thus in environments without rewards. In Figure 1, we report the results for a short, and thus more challenging, exploration horizon $(T=50)$ over Env. (i), as it is far more interpretable. Other experiments with longer horizons or over Env. (ii) can be found in Appendix C. Interestingly, at this challenging exploration regime, when looking at the joint entropy in Figure 1(a), joint and disjoint objectives perform rather well compared to mixture ones in terms of induced joint entropy, while they fail to address mixture entropy explicitly, as seen in Figure 1(b). On the other hand mixture-based objectives result in optimizing both mixture and joint entropy effectively, as one would expect by the bounds in Th. 4.1. By looking at the actual state visitation induced by the trained policies, the difference between the objectives is apparent. While optimizing joint objectives, agents exploit the high-dimensionality of the joint space to induce highly entropic distributions even without exploring the space uniformly via coordination (Fig. 1(d)); the same outcome happens in disjoint objectives, with which agents focus on over-optimizing over a restricted space loosing any incentive for coordinated exploration (Fig. 1(e)). On the other hand, mixture objectives enforce a clustering behavior (Fig. 1(e)) and result in a better efficient exploration.
Policy Pre-Training via Task-Agnostic Exploration.  More interestingly, we tested the effect of pre-training policies via task-agnostic exploration as a way to alleviate the well-known hardness of sparse-reward settings. In order to do so, we employed a multi-agent counterpart of the TRPO algorithm [36] with different pre-trained policies. First, we investigated the effect on the learning curve in the hard-exploration task of Env. (i) under long horizons ($T=150$), with a worst-case goal set on the opposite corner of the closed room. Pre-training via mixture objectives still lead to a faster learning compared to initializing the policy with a uniform distribution. On the other hand, joint objective pre-training did not lead to substantial improvements over standard initializations. More interestingly, when extremely short horizons were taken into account ($T=50$) the difference became appalling, as shown in Fig. 2(a): pre-training via mixture-based objectives leaded to faster learning and higher performances, while pre-training via disjoint objectives turned out to be even harmful (Fig. 2(b)). [^7] This is probably due to pre-training almost deterministic policies, as shown in Fig. 4 in Appendix C. Finally, we tested the zero-shot capabilities of policy pre-training on the simpler but high dimensional exploration task of Env. (ii), where the goal was sampled randomly between worst-case positions at the boundaries of the region reachable by the arm. As shown in Fig. 3(p), both joint and mixture were able to guarantee zero-shot performances via pre-training compatible with MA-TRPO after learning over $2$e4444 samples, while disjoint objectives were not. On the other hand, pre-training with joint objectives showed an extremely high-variance, leading to worst-case performances not better than the ones of random initialization. Mixture objectives on the other hand showed higher stability in guaranteeing compelling zero-shot performance.
MixtureJointDisjointRandom Initialization
(a) MA-TRPO + TRPE Pre-Training (Env. (i), $T=150$)
(b) MA-TRPO + TRPE Pre-Training (Env. (i), $T=50$)
(c) MA-TRPO + TRPE Pre-Training (Env. (ii), $T=100$)
Figure 2: Effect of pre-training in sparse-reward settings. (Left) Policies initialized with either Uniform or TRPE pre-trained policies. (Right) Policies initialized with either Zero-Mean or TRPE pre-trained policies. We report the average and 95% c.i. over 4 runs over worst-case goals.
Takeaways.  Overall, all of the experimental questions were answered: (Q4.1) Algorithm 5 is indeed able to optimize for finite-trial objectives; (Q4.2Mixture objectives enforce coordination, essential when high efficiency is required, while joint or disjoint objectives may fail to lead to relevant solutions because of under or over optimization; (Q4.3The efficient coordination through mixture objectives enforces the ability of pre-training via task-agnostic exploration to lead to faster and better training and even zero-shot generalization.
7 Conclusions and Perspectives
In this paper, we introduce a principled framework for unsupervised pre-training in MARL via task-agnostic exploration. First, we show formalize it via several different objectives, linked through performance bounds. Moreover, we theoretically characterize how the problem, even when tractable in theory, is non-trivial in practice. Then, we design a practical algorithm and we use it in a set of experiments to confirm the expected superiority of mixture objectives in practice, due to their ability to enforce efficient exploration over short horizons. Future works can build over our results in many directions, for instance by developing scalable algorithms for continuous domains.
Appendix A Related Works
Below, we summarize the most relevant work investigating related concepts.
Entropic Functionals in MARL.  A large plethora of works on both swarm robotics [21, 4] and multi-agent intrinsic motivation [among others, 10, 42, 50, 49, 41, 39] investigated the effects of employing entropic-like functions to boost exploration and performances in down-stream tasks. Importantly, these works are of empirical nature, and they do not investigate the theoretical properties of cMGs or multi-agent task-agnostic exploration, nor they propose algorithms able to pre-train policies without access to extrinsic rewards. [^8] The interested reader can refer to [25, 18] for an extensive investigation of the fundamental differences between intrinsic motivation and task-agnostic exploration. Finally, while a similar notion of cMGs was proposed in [7], their contributions are focused on the existence of equilibria and performances of centralized algorithms over infinite-trials objectives.
task-agnostic exploration and Policy Optimization.  Entropy maximization in MDPs was first introduced in [8] and then investigated extensively in various works [e.g., 24, 25, 27, 28, 18, 17, 37, 43, 38, 47, 46, 5]. Finally, we considered an IS policy gradient estimator inspired by the work of [22], yet with other forms of IS estimators, such as non-parametric k-NN estimators proposed in [25].
Appendix B Proofs of the Main Theoretical Results
In this Section, we report the full proofing steps of the Theorems and Lemmas in the main paper. In general, thought this paper, we make the following assumption for the objective function $\mathcal{F}$:
Assumption B.1 (Lipschitz).
A function $\mathcal{F}:\mathcal{A}\rightarrow\mathbb{R}$ is Lipschitz-continuous for some constant $L<\infty$, or L-Lipschitz for short, if it holds $|\mathcal{F}(x)-\mathcal{F}(y)|\leq L\|x-y\|_{1},\;\forall(x,y)\in\mathcal{A}^{2}.$
Here follow the proofs of the main results. See 4.1
Proof.
The bounds follow directly from simple yet fundamental relationships between entropies of joint, marginal and mixture distributions which can be found in [31, 12], in particular: